OpenStudy (anonymous):

Solve. log3 x – 1 = 3

OpenStudy (lgbasallote):

log_3 (x-1) = 3? change to exponential 3^3 = x-1 27 = x-1 x = 28

OpenStudy (anonymous):

thanks!!